OpenStudy (anonymous):

help!!!! Solve the following system of equations using substitution x+y=5 and x+y=-3

OpenStudy (anonymous):

I am the worst at this kind of thing.

OpenStudy (anonymous):

I will fan and medal

OpenStudy (prettygirl_shynice):

ill help

OpenStudy (anonymous):

thanks Prettygirl and Coolrabbit

OpenStudy (prettygirl_shynice):

okay the frist one x+y=5 frist step one add y to both sides x+y+−y=5+−y x=−y+5 last =y+5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1. Solve one of the equations for either "x =" or "y =". 2. Replace the "y" value in the first equation by what "y" now equals. Grab the "y" value and plug it into the other equation. 3. Solve this new equation for "x". 4. Place this new "x" value into either of the original equations in order to solve for "y". Pick the easier one to work with! 5. substitute the x and y values into both original equations. If these answers are correct, both equations will be true.

OpenStudy (prettygirl_shynice):

x+y=-3 frist Add -y to both sides.x+y+−y=−3+−y x=−y−3 last x=-y-3

OpenStudy (anonymous):

the choices are (2,3) (3,2) infinitely many No solution

OpenStudy (prettygirl_shynice):

is that what u mean by solving both are u want me to solv at the same time

OpenStudy (anonymous):

i am not sure that is the question on my homework

OpenStudy (prettygirl_shynice):

oh

OpenStudy (anonymous):

the question i posted is the one on my homework

OpenStudy (anonymous):

so what do you think it is?

OpenStudy (prettygirl_shynice):

i think u post solv e at the same time

OpenStudy (prettygirl_shynice):

it frist x+y=5for x: x+y+−y=5+−y(Add -y to both sides) x=−y+5 then Substitute−y+5forxinx+y=−3: x+y=−3 −y+5+y=−3 5=−3(Simplify both sides of the equation) 5+−5=−3+−5(Add -5 to both sides) 0=−8 and ur final answer is 0=-8

OpenStudy (prettygirl_shynice):

that the answer if u put both togeether thats it

OpenStudy (prettygirl_shynice):

is the answer on your homework